# Mixed Diffusion for 3D Indoor Scene Synthesis

### v1

@@ 5.2.2 Furniture Arrangement
5.2.3 Additional Applications
Our models can easily extend to other applications with the right choice of masking. For example, they can be applied to object type conditioned scene synthesis given a list of labels. We can even combine partial constraints on object attributes (e.g., label) with existing objects for constrained scene completion problems. We show qualitative examples in Appendix F.1.
@@ 6 Limitations
Though our method shows competitive performance in different layout generation tasks, it has several limitations: the current definition of objects as a collection of bounding box features and labels is not the most precise representation for a 3D setting, and gaining a good understanding of 3D features can potentially improve upon this aspect. In addition, MiDiffusion still requires a model retrieving strategy to compose the final 3D scene configuration. In the future we want to investigate the use of a strong 3D shape priors (e.g. in the form of a point cloud or mesh encoder) and integrating mesh and texture synthesis capabilities, similar to SceneTex [4], to enhance realism. Finally, we also plan to investigate how to remove the need for separate models per room.
@@ 7 Conclusion
In this work we introduced MiDiffusion, a novel mixed diffusion model combining DDPM and D3PM for object based 3D indoor scene synthesis. This formulation circumvents the need of adopting vector representation of semantic labels in the continuous domain as done in prior works.We designed a denoising architecture that has a time-variant transformer decoder backbone and outputs predictions separately for discrete semantic attributes and continuous geometric attributes. We proposed a unique corruption-and-masking strategy to conduct scene completion and furniture arrangement without the need to re-train our models. We provided extensive experimental evaluation using common benchmarks against state-of-the-art autoregressive and diffusion models. The results show a clear advantage of our approach over the state of the art with or without partial constraints. We also presented ablation studies supporting our design of the mixed diffusion framework as well as the denoising architecture.
@@ A.1 DDPM forward process re-parameterization, simplified loss
$L_{t-1}={\mathbb{E}}_{{\bm{x}}_{0},{\bm{\epsilon}}}\left[\frac{\beta_{t}^{2}}{2\sigma_{t}^{2}\alpha_{t}(1-\bar{\alpha}_{t})}\left\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\right\|^{2}\right]+\text{const}$ (17)
@@ A.2 D3PM reverse parameterization, auxiliary loss
Similar DDPM, [2, 16] suggested that approximating some surrogate variables gives better quality. Specifically, they trained a neural network $\tilde{p}_{\theta}(\tilde{{\bm{z}}}_{0}|{\bm{z}}_{t})$, multiplied it with the posterior $q({\bm{z}}_{t-1}|{\bm{z}}_{t},{\bm{z}}_{0})$, and marginalized $\tilde{{\bm{z}}}_{0}$ to obtain $p_{\theta}({\bm{z}}_{t-1}|{\bm{z}}_{t})$:
@@ Appendix B MiDiffusion Loss Factorization
We provide more details for the factorization step in Eq. (13). We first consider a general case comparing KL-divergence between $\tilde{q}(z)\hat{q}({\bm{x}})$ and $\tilde{p}(z)\hat{p}({\bm{x}})$ for discrete $z$ and continuous ${\bm{x}}$:
$\begin{split}D_{\mathrm{KL}}(\tilde{q}(z)\hat{q}({\bm{x}})||\tilde{p}(z)\hat{p}({\bm{x}}))=&\int_{\bm{x}}\sum_{z}\tilde{q}(z)\hat{q}({\bm{x}})\log\frac{\tilde{q}(z)\hat{q}({\bm{x}})}{\tilde{p}(z)\hat{p}({\bm{x}})}\\ =&\int_{\bm{x}}\sum_{z}\left[\tilde{q}(z)\hat{q}({\bm{x}})\log\frac{\tilde{q}(z)}{\tilde{p}(z)}+\tilde{q}(z)\hat{q}({\bm{x}})\log\frac{\hat{q}({\bm{x}})}{\hat{p}({\bm{x}})}\right]\\ =&\sum_{z}\tilde{q}(z)\log\frac{\tilde{q}(z)}{\tilde{p}(z)}\underbrace{\int_{\bm{x}}\hat{q}({\bm{x}})}_{1}+\int_{\bm{x}}\hat{q}({\bm{x}})\log\frac{\hat{q}({\bm{x}})}{\hat{p}({\bm{x}})}\underbrace{\sum_{z}\tilde{q}(z)}_{1}\\ =&D_{\mathrm{KL}}(\tilde{q}(z)||\tilde{p}(z))+D_{\mathrm{KL}}(\hat{q}({\bm{x}})||\tilde{p}({\bm{x}})).\end{split}$ (21)
@@
$\begin{split}&L_{t-1}^{mixed}\\ &=\mathbb{E}_{q(z_{t},{\bm{x}}_{t}|z_{0},{\bm{x}}_{0})}\left[D_{\mathrm{KL}}(q(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},z_{0},{\bm{x}}_{0})||p_{\theta}(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}))\right]\\ &=\mathbb{E}_{\tilde{q}(z_{t}|z_{0})\hat{q}({\bm{x}}_{t}|{\bm{x}}_{0})}\left[D_{\mathrm{KL}}(\tilde{q}(z_{t-1}|z_{t},z_{0})\hat{q}({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||\tilde{p}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}))\hat{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}))\right]\\ &=\mathbb{E}_{\tilde{q}(z_{t}|z_{0})}\left[D_{\mathrm{KL}}(\tilde{q}(z_{t-1}|z_{t},z_{0})||\tilde{p}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}})))\right]\\ &\quad+\mathbb{E}_{\hat{q}({\bm{x}}_{t}|{\bm{x}}_{0})}\left[D_{\mathrm{KL}}(\hat{q}({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||\hat{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}))\right],\end{split}$ (22)
@@ Appendix C Network and Implementation Details
We include detailed network hyper-parameters for our denoising network, and implementation for training. We plan to release our code and training configurations upon acceptance of this paper.
@@ C.1 Network Hyper-parameters
The architecture of MiDiffusion has at its core a series of 8 transformer blocks using a hidden dimension of 512 with 8 heads and feed-forward layers with an internal dimension of 2048. Following [11], we use GELU [14] as nonlinearity. The semantic attributes are embedded in a learned vector of length 512, and the geometric attributes are mapped by a 3-layer MLP, with internal dimensions of [512, 1024], to a 512-dimensional space before being fed to the transformer decoder blocks. We extract floor plan features of dimension 64 using a 4-layer PointNet [31] with internal dimensions [64, 64, 512] as in LEGO-Net [46]. The conditional input is concatenated with the learned 64-dimensional index embedding . For baselines and ablation studies, we use the default ResNet-18 [12] image feature extractor proposed by ATISS [28] and also implemented by DiffuScene [41] to compute the 64-dimensional floor plan features from binary floor plan masks. The outputs of the transformer decoder are fed to two separate MLPs to decode the semantic and geometric predictions. The semantic feature decoder is a 1-layer MLP that produces a categorical distribution over $z$, and the geometric feature decoder is a 3-layer MLP with hidden dimensions [512, 1024], producing the 8-dimensional Gaussian mean for the geometric attributes ${\bm{x}}$.
@@ C.2 Implementation
We set a fixed learning rate of $l_{r}=2e^{-4}$ using the Adam optimizer, and a dropout ratio of 0.1 for multi-head attention and feed-forward layers in the transformer blocks. We train 50k epochs on the bedroom dataset with $0.5$ learning rate decay every $10k$ epochs, and $100k$ epochs on the living and dining room datasets with $0.5$ decay every $15k$ epochs. We use a linear schedule over 1000 diffusion steps for all noise parameters in the forward process. In the discrete domain $\alpha_{t}$ and $\gamma_{t}$ range from $1-1e^{-5}$ to $0.99999$ and from $9e^{-6}$ to $0.99999$ respectively. In the continuous domain, $\beta_{t}$ starts from $1e^{-4}$ at reaches $0.02$. We train all our models on a single NVIDIA V100 GPU with under 8GB of GPU RAM usage. The training time range from around 20 hours on living and dining room datasets to about 36 hours on the bedroom dataset for a batch size of 128 scenes.
@@ Appendix D Example Synthesized Layout Images
We use texture-less object rendering on white floor plan for quantitative evaluations, since floor and object textures affect the clarity of the results. The object labels are sorted in alphabetical order for each room type, and then the associated colors are evenly sampled along a circular path using Seaborn’s [45] “hls” color palette. We render top-down layout images using simple-3dviz [1] in accordance with prior works [28, 41]. We include examples of rendered images for three comparing approaches in Fig. 789 from the floor-conditioned scene synthesis experiment in Sec 2. Some ground-truth layouts include objects slightly out of boundary. This is a problem in the raw 3D-FRONT dataset. Therefore, we inflate the room boundary by 0.1m when counting number of out of boundary objects. The bedroom datasets is the easiest (less number of objects, smaller in room size, more training data) so that all comparing approaches can generate good predictions. On the harder living room and dining room datasets, MiDiffusion clearly outperforms the baselines by generating realistic geometric arrangements with desired symmetry and alignment between objects, while respecting the floor boundary constraints. DiffuScene is also capable of generating good geometric arrangement but their architecture is less optimal for learning floor boundary constraints.
@@ E.1 Ablation Study
We provide additional geometric evaluation results for our ablation study in Sec. 5.1. Across all three dataset, there is a consistent improvement over OOB % as we modify the floor plan feature extractor and diffusion framework towards our final design of MiDiffusion. This is consistent with the evaluation metrics reported in Sec. 5.1.
Method Bedroom Dining room Living room
Obj OOB % IoU % Obj OOB % IoU % Obj OOB % IoU %
DDPM+ResNet 5.07 6.63 0.50 10.87 7.08 0.70 11.90 10.93 0.70
DDPM+PointNet 4.92 5.29 0.70 10.78 6.75 0.78 11.83 10.37 0.64
Mixed+PointNet 5.22 3.91 0.61 10.92 5.77 0.91 12.28 8.71 0.69
Ground truth 5.22 3.37 0.24 11.11 0.73 0.48 11.67 1.55 0.27
E.2 Unconditional 3D Scene Synthesis
DiffuScene was originally proposed for 3D scene synthesis without floor conditioning. Although this is not the main focus of this paper, we provide experimental results comparing MiDiffusion and DiffuScene for completeness using the released model weights. DiffuScene used squares of the same size to render the layouts, which cuts off some objects in the dining and living rooms. We therefore re-compute DiffuScene results using their released models and render layouts to dataset specific sizes. The results are summarized in Table 8. Our approach works better than DiffuScene on the dining room dataset, comparably on the living room dataset, and slightly worse on the bedroom dataset.
Table 9 shows results using DiffuScene’s default rendering and evaluation approach over the same set of results as Table 8. Specifically, we render all ground-truth and synthesized layouts to a 3.1m square. We compare the synthetic results with the ground-truth training set in accordance with DiffuScene’s evaluation script. We include the original results in  [41] in the first row as a reference. Using pre-trained models and open-sourced evaluation scripts, we are able generate DiffuScene results close to publication values. The only exception is CA%, for which DiffuScene does not release their classification network. Therefore, we use the AlexNet [22] based classifier by ATISS [28]. These results are consistent with Table 8 that MiDiffusion behaves sub-optimally on the bedroom dataset. Our models are quite similar in performance compared to DiffuScene on dining room and living room datasets, even though they are not specifically designed for this task.

### v2

@@ 5.2 Floor Plan Conditioned 3D Scene Synthesis with Object Constraints
Recall from Sec. 4.3 that we can re-use the same MiDiffusion models for problems involving partial object constraints through a corruption-and-masking strategy. This approach forces the relevant object attributes to follow the reverse of a pre-computed corruption process in the denoising step, without the need of task-specific retraining. Although it is also feasible to directly mask out relevant features by known inputs in each denoising step as proposed in [27], this approach creates out-of-distribution latent variables in the intermediate denoising step and is therefore less principled. We show comparison between the two masking approaches and additional applications in Appendix.
5.2.1 Scene completion
Figure 3: Scene completion. Bedroom (top) and living room (bottom) scene completion examples.
We first compare our approach with baselines for scene completion (i.e. scene synthesis conditioned on existing objects). We train the DiffuScene-SC variant, which includes an additional module to learn existing object features as a conditional input, using their suggested hyper-parameters and providing 3 existing objects during training. The number of existing objects can be arbitrary at test time. We use the same models for ATISS and MiDiffusion. We use the proposed corruption-and-masking strategy over MiDiffusion models trained in Section 5.1. Fig. 3 shows a simple example of completing a bedroom design that originally only consists of a single bed and another more complex example for completion of a living space. Note that MiDiffusion is able to generate a natural and symmetric set of furniture layout.
We show quantitative evaluations in Table 5. This is an easier problem for ATISS and DiffuScene than the previous problem. In particular, ATISS is autoregressive and therefore existing objects allow the models to skip the first few iterations reducing the chance for mistakes. DiffuScene-SC is, on the other hand, specifically designed and trained for this task. Compared to results in Table 1, these approaches indeed benefit from using existing objects as anchors. Nevertheless, without any retraining, our approach still outperforms these baselines when given only 1 existing object. When given 3 objects, our models are still better on the dining room and living room datasets, however, are outperformed by DiffuScene-SC on the bedroom dataset. Note that DiffuScene-SC is particularly optimized for the 3 objects scenario and that the bedroom data has only an average of only 5.22 objects. Hence, specifically trained networks, such as DiffuScene-SC, might be better suited given the very high percentage of pre-existing objects.
5.2.2 Furniture Arrangement
Figure 4: Furniture arrangement. Different table-chair arrangements (top) and different furniture placement directions (bottom).
MiDiffusion can also be used for furniture arrangement, a scene synthesis problem conditioned on object labels and bounding box sizes without an initial configuration. We use our proposed corruption-and-masking strategy over these attributes to test our models in this setting, and compare the results against DiffuScene’s furniture arrangement variant (DiffuScene-FA). ATISS models are not suited for this task, since replacing the predicted object label by constraints breaks the stopping condition in their autoregressive pipeline. Fig. 4 shows two pairs of example results with each pair generated from the same input. We show quantitative results in Table 6. Without re-training, our models outperform DiffuScene-FA in all metrics.
@@ 6 Limitations
Though our method shows competitive performance in different layout generation tasks, it has several limitations: the current definition of objects as a collection of bounding box features and labels is not the most precise representation for a 3D setting, and gaining a good understanding of 3D features can potentially improve upon this aspect. In addition, MiDiffusion still requires a model retrieving strategy to compose the final 3D scene configuration. In the future we want to investigate the use of strong 3D shape priors (in the form of a point cloud or mesh encoder) and integrating mesh and texture synthesis capabilities such as SceneTex [4], to enhance realism.
@@ 7 Conclusion
In this work we introduced MiDiffusion, a novel mixed diffusion model combining DDPM and D3PM for object based 3D indoor scene synthesis. This formulation circumvents the need of adopting one-hot representations of semantic labels in the continuous domain as done in prior works.We designed a denoising architecture that has a time-variant transformer decoder backbone and outputs predictions separately for discrete semantic attributes and continuous geometric attributes. We proposed a corruption-and-masking strategy to conduct scene completion and furniture arrangement without the need to re-train our models. We provided extensive experimental evaluation using common benchmarks against state-of-the-art autoregressive and diffusion models. The results show a clear advantage of our approach over the state-of-the-art with or without partial constraints. We also presented ablation studies supporting our design of the mixed diffusion framework as well as the denoising architecture.
@@ A.1 DDPM forward process re-parameterization, simplified loss
$\begin{split}L_{t-1}=&{\mathbb{E}}_{{\bm{x}}_{0},{\bm{\epsilon}}}\left[\frac{\beta_{t}^{2}}{2\sigma_{t}^{2}\alpha_{t}(1-\bar{\alpha}_{t})}\right.\\ &\hskip 28.45274pt\left.\left\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\right\|^{2}\right]\\ &+\text{const}\end{split}$ (18)
@@ A.2 D3PM reverse parameterization, auxiliary loss
Similar to DDPM, [2, 19] suggested that approximating some surrogate variables gives better quality. Specifically, they trained a neural network $\tilde{p}_{\theta}(\tilde{{\bm{z}}}_{0}|{\bm{z}}_{t})$, multiplied it with the posterior $q({\bm{z}}_{t-1}|{\bm{z}}_{t},{\bm{z}}_{0})$, and marginalized $\tilde{{\bm{z}}}_{0}$ to obtain $p_{\theta}({\bm{z}}_{t-1}|{\bm{z}}_{t})$:
@@ Appendix B MiDiffusion Loss Factorization
We provide more details for the factorization step in Eq. (14). We first consider a general case comparing the KL-divergence between $\tilde{q}(z)\hat{q}({\bm{x}})$ and $\tilde{p}(z)\hat{p}({\bm{x}})$ for discrete $z$ and continuous ${\bm{x}}$:
$\begin{split}&D_{\mathrm{KL}}(\tilde{q}(z)\hat{q}({\bm{x}})||\tilde{p}(z)\hat{p}({\bm{x}}))=\int_{\bm{x}}\sum_{z}\tilde{q}(z)\hat{q}({\bm{x}})\log\frac{\tilde{q}(z)\hat{q}({\bm{x}})}{\tilde{p}(z)\hat{p}({\bm{x}})}\\ &=\int_{\bm{x}}\sum_{z}\left[\tilde{q}(z)\hat{q}({\bm{x}})\log\frac{\tilde{q}(z)}{\tilde{p}(z)}+\tilde{q}(z)\hat{q}({\bm{x}})\log\frac{\hat{q}({\bm{x}})}{\hat{p}({\bm{x}})}\right]\\ &=\sum_{z}\tilde{q}(z)\log\frac{\tilde{q}(z)}{\tilde{p}(z)}\underbrace{\int_{\bm{x}}\hat{q}({\bm{x}})}_{1}+\int_{\bm{x}}\hat{q}({\bm{x}})\log\frac{\hat{q}({\bm{x}})}{\hat{p}({\bm{x}})}\underbrace{\sum_{z}\tilde{q}(z)}_{1}\\ &=D_{\mathrm{KL}}(\tilde{q}(z)||\tilde{p}(z))+D_{\mathrm{KL}}(\hat{q}({\bm{x}})||\tilde{p}({\bm{x}})).\end{split}$ (22)
@@
$\begin{split}&L_{t-1}^{mixed}=\mathbb{E}_{q(z_{t},{\bm{x}}_{t}|z_{0},{\bm{x}}_{0})}\left[D_{\mathrm{KL}}(q(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},z_{0},{\bm{x}}_{0})\right.\\ &\hskip 128.0374pt\left.||p_{\theta}(z_{t-1},{\bm{x}}_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}))\right]\\ &=\mathbb{E}_{\tilde{q}(z_{t}|z_{0})\hat{q}({\bm{x}}_{t}|{\bm{x}}_{0})}\left[D_{\mathrm{KL}}(\tilde{q}(z_{t-1}|z_{t},z_{0})\hat{q}({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})\right.\\ &\hskip 85.35826pt\left.||\tilde{p}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}}))\hat{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}))\right]\\ &=\mathbb{E}_{\tilde{q}(z_{t}|z_{0})}\left[D_{\mathrm{KL}}(\tilde{q}(z_{t-1}|z_{t},z_{0})||\tilde{p}_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}})))\right]\\ &\quad+\mathbb{E}_{\hat{q}({\bm{x}}_{t}|{\bm{x}}_{0})}\left[D_{\mathrm{KL}}(\hat{q}({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})||\hat{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}}))\right],\end{split}$ (23)
@@ Appendix C Network and Implementation Details
We include detailed network hyper-parameters for our denoising network, and implementation for training.
@@ C.1 Network Hyper-parameters
The architecture of MiDiffusion has at its core a series of 8 transformer blocks using a hidden dimension of 512 with 4 heads and feed-forward layers with an internal dimension of 2048. Following [13], we use GELU [16] as nonlinearity. The semantic attributes are embedded in a learned vector of length 512, and the geometric attributes are mapped by a 3-layer MLP, with internal dimensions of [512, 1024], to a 512-dimensional space before being fed to the transformer decoder blocks. We extract floor plan features of dimension 64 using a 4-layer PointNet [35] with internal dimensions [64, 64, 512] as in LEGO-Net [51]. For baselines and ablation studies, we use the default ResNet-18 [14] image feature extractor proposed by ATISS [32] and also implemented by DiffuScene [46] to compute the 64-dimensional floor plan features from binary floor plan masks. The outputs of the transformer decoder are fed to two separate MLPs to decode the semantic and geometric predictions. The semantic feature decoder is a 1-layer MLP that produces a categorical distribution over $z$, and the geometric feature decoder is a 3-layer MLP with hidden dimensions [512, 1024], producing the 8-dimensional Gaussian mean for the geometric attributes ${\bm{x}}$.
@@ C.2 Implementation
We set a fixed learning rate of $l_{r}=2e^{-4}$ using the Adam optimizer, and a dropout ratio of 0.1 for multi-head attention and feed-forward layers in the transformer blocks. We train 50k epochs on the bedroom dataset with $0.5$ learning rate decay every $10k$ epochs, and $100k$ epochs on the living and dining room datasets with $0.5$ decay every $15k$ epochs. We use a linear schedule over 1000 diffusion steps for all noise parameters in the forward process. In the discrete domain $\alpha_{t}$ and $\gamma_{t}$ range from $1-1e^{-5}$ to $0.99999$ and from $9e^{-6}$ to $0.99999$ respectively. In the continuous domain, $\beta_{t}$ starts from $1e^{-4}$ at reaches $0.02$. We train all our models on a single NVIDIA V100 GPU with under 8GB of GPU RAM usage. The training time range from around 20 hours on living and dining room datasets to about 32 hours on the bedroom dataset for a batch size of 512 scenes.
Method Bedroom Dining room Living room
2-stage + PointNet 65.54 1.53 57.83 $\pm$ 4.89 3.50 38.80 7.02 61.53 $\pm$ 7.85 5.08 35.35 6.30 61.97 $\pm$ 6.55 4.73
2-stage (gt labels) + PointNet 64.37 0.67 56.19 $\pm$ 5.45 32.47 5.07 56.29 $\pm$ 4.80 32.26 5.86 58.85 $\pm$ 6.29
Mixed+PointNet 63.71 0.24 53.65 $\pm$ 2.46 1.05 30.63 2.74 52.10 $\pm$ 1.72 1.90 28.60 1.64 54.04 $\pm$ 3.33 1.54
Table 7: Additional ablation study for floor-conditioned 3D scene synthesis.
Method Bedroom Dining room Living room
Obj OOB % IoU % Obj OOB % IoU % Obj OOB % IoU %
2stage + PointNet 5.04 5.67 0.72 11.30 11.80 1.57 11.50 13.35 1.48
2stage (gt labels) + PointNet 5.20 0.61 8.87 1.18 11.80 1.28
DDPM+ResNet 5.07 6.63 0.50 10.87 7.08 0.70 11.90 10.93 0.70
Mixed+ResNet 5.15 4.43 0.44 10.78 7.09 0.73 11.91 10.60 0.70
DDPM+PointNet 4.92 5.29 0.70 10.78 6.75 0.78 11.83 10.37 0.64
Mixed+PointNet 5.22 3.91 0.61 10.92 5.77 0.91 11.91 7.93 0.98
Ground truth 5.22 3.37 0.24 11.11 0.73 0.48 11.67 1.55 0.27
Method Bedroom Dining room Living room
MiDiffusion + DM 59.43 -0.07 7.71 0.56 32.17 4.46 6.49 0.95 30.47 3.65 8.72 0.84
Table 9: Additional evaluation results for furniture arrangement experiment.
@@ Appendix D Example Synthesized Layout Images
We use texture-less objects rendered on white floor plans for quantitative evaluations, since floor and object textures affect the clarity of the results. The object labels are sorted in alphabetical order for each room type, and then the associated colors are evenly sampled along a circular path using Seaborn’s [50] “hls” color palette. We render top-down layout images using simple-3dviz [1] in accordance with prior works [32, 46]. We include examples of rendered images for three comparing approaches in Fig. 567 from the floor-conditioned scene synthesis experiment in Sec 2. Some ground-truth layouts include objects slightly out of boundary. This is a problem in the raw 3D-FRONT dataset. Therefore, we inflate the room boundary by 0.1m when counting the number of out of boundary objects. The bedroom datasets is the easiest (less number of objects, smaller in room size, more training data) so that all comparing approaches can generate good predictions. On the harder living room and dining room datasets, MiDiffusion clearly outperforms the baselines by generating realistic geometric arrangements with desired symmetry and alignment between objects, while respecting the floor boundary constraints. DiffuScene is also capable of generating good geometric arrangements but their architecture is less optimal for learning floor boundary constraints.
@@ Appendix E Additional Experimental Evaluations
We provide supplementary results to complement Sec. 5. We begin with additional results from the ablation studies to further support our proposed mixed diffusion formulation and network design. Then, we present more results on scene synthesis with partial constraints, including a comparison with the direct masking strategy and applications to other conditional setups. Lastly, we show quantitative comparison between MiDiffusion and DiffuScene in the unconditioned scene synthesis experiment, which is original target application of DiffuScene.
@@ E.1 Ablation Study
We show an additional ablation study comparing MiDiffusion (also referred as Mixed+PointNet in ablation studies) and the 2-stage approach to supplement Table 3. The second row of Table 7 reports results of a 2-stage variant where ground-truth semantic labels are sent to the geometric diffusion stage. While bypassing the first semantic diffusion stage improves performance, this variant still under-performs compared to MiDiffusion. This suggests that even with a perfect first-stage label predictor, the 2-stage approach still produces less realistic scene layouts compared to ours.
We show object evaluation results in Table 8. Across all three datasets, there is a consistent improvement over OOB % as we modify the floor plan feature extractor and diffusion framework towards our final design of MiDiffusion. This is consistent with the main results in Table 3.
E.2 Masking Comparison
In Sec. 5.2, we show that, with the proposed corruption-and-masking strategy, MiDiffusion models are able to generate room layouts under various types of object constraints. Compared to ours, direct masking at each denoising step is less principled. We provide an example comparison in Table 9 to demonstrate that our proposed masking strategy yields better results.
E.3 Additional Applications
Our models can easily extend to other applications with different masking ranges. For example, they can be applied to object category constrained scene synthesis given a set of desired object labels. Fig. 8 shows two examples of label constrained scene synthesis. In particular, the dining room scene is conditioned on five dining chairs, which occur less often than an even number of chairs. Our model is still able to generate reasonable layouts. We can even combine partial constraints on object attributes (e.g. label) with existing objects. Fig. 9 shows examples of such label constrained scene completion problems.
Figure 8: Label constrained scene synthesis. Bedroom (left) and dining room (bottom) examples.
Figure 9: Scene completion with label constraints. Bedroom (top) and living room (bottom) label constrained scene completion examples.
E.4 Unconditional 3D Scene Synthesis